# Dynamic planning in hierarchical active inference

### v2

@@ 2.3 Intention modulation and multi-step behavior
Figure 6: (a) In this task, the agent has to first touch a moving target (red circle) and then reach the home position (grey square). Additionally, both elements have to be inferred through sensory observations before the agent starts moving. This phase of pure perceptual inference is realized by setting both hidden causes to 00. The second transition is done by a tactile belief. Note how even during the second movement, the agent continues to track the belief over the target angle. See [43, 75] for more details. (b) The first graph shows the evolution of the real arm angle $\bm{x}_{a}$ and its belief $\bm{\mu}_{a}$, the target angle $\bm{x}_{t}$ and its belief $\bm{\mu}_{t}$, the home angle $\bm{x}_{h}$ and its belief $\bm{\mu}_{h}$. The second graph shows the evolution of the hidden causes associated with the two intentions, $\nu_{reach}$ and $\nu_{return}$. The third graph shows the evolution of the intention prediction errors $\bm{e}_{i,reach}$ and $\bm{e}_{i,return}$, and the dynamics functions $\bm{f}_{reach}$ and $\bm{f}_{return}$. The last graph shows the evolution of the tactile observation $\bm{o}_{t}$, and its belief $\bm{\mu}_{t}$.
@@ 3 Hierarchical models
So far, we have introduced several units with two kinds of inputs – a prior over the hidden states and a prior over the hidden causes – and one kind of output – the 0th-order observation. In this chapter, we focus on how to combine such units in a single network to achieve a more advanced and efficient control. For this, we will make use of the first input, leaving the discussion about the second to the next chapter.
@@ 3.1 Intrinsic and extrinsic causes
The last unit presented affords a multi-step behavior in continuous time that accounts, to some extent, for dynamic elements of the environment. However, in all the previous simulations we just considered a single-DoF arm, while in real-life applications we generally deal with much more complex kinematic structures such as the human body. In this case, there is no more a one-to-one mapping between joint angles and Cartesian positions, so we need to distinguish between proprioceptive and exteroceptive (e.g., visual) observations. As in optimal control, continuous-time active inference considers three reference frames and two inversions: an extrinsic signal (e.g., encoding the Cartesian position of a target) is first transformed in an intrinsic signal (e.g., encoding the joint angles configuration corresponding to the hand at the target) through inverse kinematics, which is in turn converted to the actual motor control signals (e.g., joint torques) through inverse dynamics [94]. These two processes are also attributed to the human brain [95, 96, 97], but there is a substantial difference between optimal control and active inference regarding how they unfold in practice. As mentioned in the previous chapter, in active inference the motor commands are replaced by proprioceptive prediction errors that are suppressed through spinal reflex arcs [34]. As a consequence, inverse dynamics becomes easier because action is put aside and the agent has just to know the mapping from proprioceptive states to motor commands – see Equation 9.
But what about inverse kinematics? Recall the perspective that we mentioned in the previous chapter, i.e., that “it is an object I want to reach that generates my movements”. Turning optimal control upside down, active inference posits that action is driven by the proprioceptive consequences (e.g., changes in limb lengths) of extrinsic causes (e.g., a target) [33]. Intuitively, one could model an extrinsic movement as in Figure 7(a), i.e., with the following dynamics and likelihood functions:
@@
where $\bm{T}$ is the forward kinematics returning the hand position, and $\bm{J}$ is the Jacobian matrix.
Figure 7: (a) Factor graph of an active inference model commonly used for kinematics. Hidden causes of a single level represent a target to reach, while hidden states define the joint angles of the kinematic chain, generating proprioceptive predictions through $\bm{g}_{p}$ and visual predictions through $\bm{g}_{v}$. Note that forward and inverse kinematics are duplicated, also requiring the likelihood $\bm{g}_{v}$ to be embedded into the dynamics function. (b) Factor graph of an alternative hierarchical model for kinematics. Two different units (left and right blocks) encode information about arm joint angles and hand position, respectively generating proprioceptive and visual predictions. The two levels are linked by the likelihood function $\bm{g}_{e}$ performing forward kinematics. Inverse kinematics and goal-directed behavior arise naturally through inference; moreover, both levels can express their own dynamics, affording more advanced control.
In short, the hand – expressed in terms of joint angles of the whole arm – is embedded in the hidden states, while the target to reach – expressed as a Cartesian position – is encoded in the hidden causes. The proprioceptive states needed for movement are found by first using an inverse kinematic model directly as a forward model into the dynamics function, e.g., through a Jacobian transpose or a pseudoinverse [85, 86, 87, 37, 28, 98, 93]; and then, by generating a prediction through the proprioceptive likelihood $\bm{g}_{p}$ (which here is a simple identity mapping). The visual likelihood $\bm{g}_{v}$ is finally used to update the target location and further refine the inference of the agent’s configuration. This approach implies that an extrinsic reference frame is inverted to generate an intrinsic state, which is in turn transformed again in the first domain to be compared with visual observations. As a result, forward and inverse kinematics are performed twice, once in the dynamics function and once in the forward and backward passes of perceptual inference, when propagating the visual prediction error $\bm{\varepsilon}_{v}$:
$\partial\bm{g}_{v}^{T}\bm{\varepsilon}_{v}=\begin{bmatrix}\bm{J}\\ \bm{1}\end{bmatrix}(\bm{o}_{v}-\begin{bmatrix}\bm{T}(\bm{x})&\bm{v}\end{bmatrix})$ (23)
If the predictions are not temporarily stored, this requires increased computational demand and memory. Crucially, there is an additional issue regarding biological plausibility: using sensory-level attractors within the dynamics function means that a unit is aware of and can use intra-level part of the likelihood prediction – which is generally assumed to go all down to the sensorium – and its inverse mapping, which are lower-level features. Finally, the model in Figure 7(a) does not let the agent express paths in extrinsic coordinates needed, e.g., for realizing linear or circular motions, or for imposing constraints in both intrinsic and extrinsic domains such as when walking with a glass in hand.
@@
We can instead exploit Equation 23 and follow the natural flow of the generative process to avoid duplicated computations, as displayed in Figure 7(b). This model relies on two hierarchical levels, where an intrinsic unit (encoding the arm joint angles) is placed at the top and generates predictions through forward kinematics for an extrinsic unit (encoding the Cartesian position of the target) [44]:
$\bm{x}_{e}=\bm{g}(\bm{x}_{i})=\bm{T}(\bm{x}_{i})+\bm{w}_{i}$ (24)
@@
The goal-directed behavior of Equation 22 arises naturally via backpropagation of the extrinsic prediction error $\bm{\varepsilon}_{e}$:
$\partial\bm{g}_{e}^{T}\bm{\varepsilon}_{e}=\bm{J}^{T}(\bm{x}_{e}-\bm{T}(\bm{x}_{i}))$ (25)
@@
Having a complete unit that deals with extrinsic information – which is thus not embedded into the hidden causes of the intrinsic unit – allows the agent to specify their dynamics, leading to an efficient decomposition between intrinsic and extrinsic attractors, and between proprioceptive and visual observations – as exemplified in the simulations of Figure 8. Note the similarity of Equation 25 with Equations 22 and 23: if in the model of Figure 7(a) we had two different forward and inverse kinematics either for goal-directed behavior or for predicting current observations, in this case what is compared with the observations already contains a bias toward preferred states, without the need for sensory-level attractors within the dynamics function.
@@
Although the generative model follows the forward flow of optimal control, the relationship between proprioceptive consequences and extrinsic causes peculiar to active inference still holds because the kinematic inversion regards a high-level process that manipulates abstract (intrinsic or extrinsic) representations, and both of them concur to generate low-level proprioceptive states. As noted by Adams and colleagues, “The key distinction is not about mapping from desired states in an extrinsic (kinematic) frame to an intrinsic (dynamic) frame of reference, but the mapping from desired states (in either frame) to motor commands[34]. Having said this, there is a significant difference between the two models represented in Figure 7, which can be compared to the two supervised learning modes of predictive coding [16]: a forward mode that fixes the latent states to the labels and the observations to the data can generate highly accurate images of digits, while the inverse classification task is more difficult as there is no univocal mapping between labels and data; instead, a backward mode that fixes the latent states to the data and the observations to the labels achieves high performances on classification but falls short when generating images. Based on this, we can interpret the model of Figure 7(a) as a backward mode that would rapidly generate a proper kinematic configuration with the hand at the target, but that would hardly infer from proprioception the hand position needed to plan movements. Conversely, we can interpret the model of Figure 7(b) as a forward mode that would generate with high accuracy the hand position, but that would find it difficult to infer the kinematic configuration needed to actually realize movement.
@@ 3.2 A module for iterative transformations
The model in Figure 7(b) introduced a hierarchical dependency between two (intrinsic and extrinsic) levels, made possible by a connection between hidden states. Instead, the typical approach in continuous-time active inference is to let the hidden states and causes of a level exchange information with the hidden causes (and not the hidden states) of the subordinate level, as shown in Figure 9(a). While this allows one to impose a dynamic trajectory for the unit below, specifying fixed setpoints to the 0th-order hidden states is not as straightforward, since the dynamics prediction error generated from the hidden causes would have to travel back to the previous temporal orders. As clear from Figure 7(b), a connection between hidden states is of extreme importance when designing hierarchical models. In fact – as represented in Figure 9(b) – it is fundamental in defining the initial state of slower temporal scales in discrete models, e.g., in pictographic reading [48]. An analogy could be also made regarding the hierarchical connectivity of PCNs, as units are connected in a multiple-input and multiple-output system, defining static priors for the subordinate levels [16] – as shown in Figure 9(c).
@@
Figure 9: (a) Factor graph of the most common choice of connections between two continuous levels. Hidden states $\bm{x}^{(i)}$ and hidden causes $\bm{v}^{(i)}$ of level $i$ generate – through the likelihood function $\bm{g}^{(i)}$ – the hidden causes $\bm{v}^{(i+1)}$ of level $i+1$. (b) Connections between two discrete levels. Hidden states $\bm{s}_{1}^{(i)}$ of level $i$ generate – through the prior matrix $\bm{D}^{(i)}$ – the hidden states of the present $\bm{s}_{1}^{(i+1)}$ of level $i+1$. We will cover discrete models in the next chapter. (c) Connections between two levels in PCNs. Hidden causes $\bm{v}^{(i)}$ of level $i$ receive an average of prediction errors from units at level $i-1$, and impose a prior over units at level $i+1$. Typically, the likelihood function $\bm{g}^{(i)}$ performs a combination of independent units, passed to a nonlinear activation function. Standard PCNs only permit representing the causal structure of a system, without modeling internal dynamics. (d) Factor graph of a level with multiple inputs and outputs from independent units, with a similar connectivity of PCNs plus model dynamics. The observation $\bm{o}^{(i,j)}$ becomes the 0th-order hidden state $\bm{x}^{(i+1,j)}$ of the level below, while the prior $\bm{\eta}_{x}^{(i,j)}$ becomes the 0th-order hidden state $\bm{x}^{(i-1,j)}$ of the level above. (e) A network of IE modules. A signal expressed in an extrinsic reference frame $\bm{x}_{e}^{(i-1)}$, along with an intrinsic signal $\bm{x}_{i}^{(i,j)}$ (e.g., angle for rotation or length for translation), is passed to a function $\bm{g}_{e}^{(i,j)}$, generating a new extrinsic signal $\bm{x}_{e}^{(i,j)}$.
@@
Following these two examples and the previous kinematic model, we use the observation of a level to directly bias the 0th-order hidden states of the level below. As a result, the observation prediction error $\bm{\varepsilon}_{o}$ and the prior prediction error $\bm{\varepsilon}_{\eta_{x}}$ of Equation 6 is expressed by the same variable:
$\bm{\varepsilon}_{o}^{(i)}=\bm{\mu}^{(i+1)}-\bm{g}^{(i)}(\bm{\mu}^{(i)})$ (26)
@@
where the hierarchical level is indicated with a superscript and lower levels are denoted by increasing numbers. We can then design a multiple-input and multiple-output system wherein a level imposes and receives priors and observations to several units, as in Figure 9(d). The computation of the free energy in Equation 5 remains unchanged, and the update of the hidden states turns into the following:
$\dot{\tilde{\bm{\mu}}}^{(i,j)}=\begin{bmatrix}\bm{\mu}^{\prime{(i,j)}}-\sum_{k}\bm{\pi}_{o}^{(i-1,k)}\bm{\varepsilon}_{o}^{(i-1,k)}+\sum_{l}\partial_{\mu^{(i,j)}}\bm{g}^{(i,l)T}\bm{\pi}_{o}^{(i,l)}\bm{\varepsilon}_{o}^{(i,l)}+\partial_{\mu^{(i,j)}}\bm{f}^{(i,j)T}\bm{\pi}^{(i,j)}_{x}\bm{\varepsilon}^{(i,j)}_{x}\\ \\ -\bm{\pi}^{(i,j)}_{x}\bm{\varepsilon}^{(i,j)}_{x}\end{bmatrix}$ (27)
where the superscript notation $(i,j)$ indicates the $i$th hierarchical level and the $j$th element within the same level. As evident, this is a similar hierarchical connectivity of PCNs – wherein we highlighted an average of forward and backward prediction errors from independent units – but with the addition of model dynamics represented by the leftmost and rightmost terms of the equation.
@@
What advantages do deep hierarchical models carry compared to a shallow agent? If we consider the structure of Figure 7(b), although affording a more advanced control with respect to the model in Figure 7(a), its uses are still limited to solving simple tasks, e.g., performing operations with the hand. While simultaneous coordination of multiple limbs is certainly possible, it would require complex dynamics functions, with complexity increasing with the number of joints and ramifications of the kinematic chain. Critically, a shallow agent would not be capable of capturing the hierarchical causal relationships inherent to the generative process, allowing one to predict and anticipate the local exchange of forces that would unfold whenever a biased belief over bodily states produces a movement. As mentioned in the Introduction, a deep model is also required if one has to flexibly use external tools for manipulation tasks. Besides the roto-translations occurring in forward kinematics, iterative transformations are also essential in computer vision – where an image can be subject to scaling, shearing, or projection – and, more in general, whenever changing basis of a coordinate vector.
@@
For these reasons, we can generalize the last model and construct an Intrinsic-Extrinsic (or IE) module [44, 58, 99]. This module is composed of two units and its role is to perform iterative transformations between reference frames. In brief, a unit $\mathcal{U}_{e}^{(i-1)}$ encodes a signal in an extrinsic reference frame, while a second unit $\mathcal{U}_{i}^{(i)}$ contains a generic intrinsic transformation. Applying the latter to the first signal returns a new extrinsic reference frame embedded in a unit $\mathcal{U}_{e}^{(i)}$. More formally, we can define a likelihood function $\bm{g}_{e}$ such that:
$\bm{x}_{e}^{(i)}=\bm{g}_{e}^{(i)}(\bm{x}_{i}^{(i)},\bm{x}_{e}^{(i-1)})=\bm{T}^{(i)}(\bm{x}_{i}^{(i)})\cdot\bm{x}_{e}^{(i-1)}+\bm{w}_{e}^{(i)}$ (28)

### v3

@@ 2.3 Intention modulation and object affordances
Figure 6: (a) In this task, the agent has to first touch the moving ball (in red) and then track the moving square (in grey). The transition is done by a tactile belief. Note how throughout the task, the agent maintains potential body configurations related to the two objects. See [64, 66] for more details. (b) The first graph shows the difference between the estimated and real hand positions (dotted blue), the difference between the potential hand position related to the ball and the actual ball position (dotted red), the difference between the potential hand position related to the square and the actual square position (dotted green), the difference between the real hand and ball positions (solid dark red), and between the real hand and square positions (solid dark green). The second graph shows the evolution of the hidden causes associated with the two intentions, $\mu_{v,ball}$ and $\mu_{v,square}$. The third graph shows the evolution of the attractors $\bm{e}_{i,ball}$ and $\bm{e}_{i,square}$, and the dynamics functions $\bm{f}_{ball}$ and $\bm{f}_{square}$, regarding the first joint angle. The last graph shows the evolution of the tactile observation $\bm{o}_{t}$, and its belief $\bm{\mu}_{t}$.
The generative model is shown in Figure 5(b). The update rule for the 0th-order hidden states becomes:
$\dot{\bm{\mu}}_{x}=\bm{\mu}_{x}^{\prime}-\bm{\Pi}_{\eta,x}\bm{\varepsilon}_{\eta,x}+\partial_{x}\bm{g}_{p}^{T}\bm{\Pi}_{o,p}\bm{\varepsilon}_{o,p}+\sum_{n}^{N}\partial_{x}\bm{g}_{v}^{T}\bm{\Pi}_{o,v,n}\bm{\varepsilon}_{o,v,n}+\partial_{x}\bm{f}^{T}\bm{\Pi}_{x}\bm{\varepsilon}_{x}$ (16)
In particular, the dynamics prediction error:
$\bm{\varepsilon}_{x}=\bm{\mu}_{x}^{\prime}-\bm{\eta}_{x}^{\prime}$ (17)
realizes an average trajectory that the agent predicts for the current situation. This approach is effective for two reasons. First, it allows defining a composite movement in terms of simpler subgoals, which can be tackled separately; this can be helpful, for instance, if one has to analyze the behavior of an agent when subject to two or more opposing priors [64]. Second, a simple multi-step behavior without planning can be achieved [66], wherein one just needs to modulate the hidden causes. Transitions between continuous trajectories could then be realized by higher-level priors, e.g., a belief of tactile sensations. Third, and most importantly, it permits maintaining in parallel potential body configurations related to the objects to be manipulated – thus providing efficient transitions between movements – and encoding the objects according to their affordances and the agent’s intentions (e.g., grasping a cup by the handle or with the whole hand). These features are illustrated in the simulation of Figure 6, showing a two-step reaching task with moving objects.
@@ 3 Hierarchical models
So far, we have discussed several units with two kinds of inputs – a prior over the hidden states and a prior over the hidden causes – and one kind of output – the 0th-order observation. In this chapter, we focus on combining such units in a single network to achieve a more advanced and efficient control. For this, we will make use of the first input, leaving the discussion about the second to the next chapter.
In hierarchical active inference models, units are arranged in layers so that the output of one layer provides input to the subordinate layers. This architecture permits representing complex data, such as convolutions models or nonlinear time series [57]. In motor control, a (deep) hierarchy of goals integrates control and motivational streams of the brain [82]. For robotics, a hierarchical kinematic model can be designed in continuous time, wherein each unit encodes a certain Degree of Freedom (DoF) in intrinsic and extrinsic reference frames [80]. This permits realizing advanced movements that involve simultaneous coordination of multiple limbs, e.g., moving with a glass in hand. This hierarchical structure can be generalized to perform homogeneous transformations between reference frames, e.g., perspective projections [83].
@@ 3.1 Intrinsic and extrinsic causes
The last unit presented affords a multi-step behavior in continuous time that accounts for object affordances and, to some extent, for dynamic elements of the environment. However, it can only estimate body configurations, while in real-life applications we constantly plan movements in the spatial domain. Further, it only generates visual predictions of an object related to a single DoF (e.g., the hand), while we generally deal with much more complex kinematic structures with different branches such as the human body. As in optimal control, continuous-time active inference considers three reference frames and two inversions: an extrinsic signal (e.g., encoding the Cartesian position of a target) is first transformed in an intrinsic signal (e.g., encoding the joint angles configuration corresponding to the hand at the target) through inverse kinematics, which is in turn converted to the actual motor control signals (e.g., joint torques) through inverse dynamics [84]. These two processes are also attributed to the human brain [85, 86, 87], but there is a substantial difference between optimal control and active inference regarding how they unfold in practice. As mentioned in the previous chapter, in active inference the motor commands are replaced by proprioceptive prediction errors that are suppressed through spinal reflex arcs [38]. As a consequence, inverse dynamics becomes easier because action is put aside and the agent has just to know the mapping from proprioceptive states to motor commands – see Equation 7.
But what about inverse kinematics? Recall the perspective that we mentioned in the previous chapter, i.e., that “it is an object I want to reach that generates my movements”. Turning optimal control upside down, active inference posits that action is driven by the proprioceptive consequences (e.g., changes in muscle lengths) of extrinsic causes (e.g., a limb position) [37]. Intuitively, one could model an extrinsic movement as in Figure 7(a), i.e., with the following dynamics and likelihood functions:
@@
where $\bm{x}$ are the arm joint angles, $\bm{v}$ is the target position to reach, $\bm{T}$ is the forward kinematics returning the hand position, and $\bm{J}$ is its Jacobian matrix.
Figure 7: (a) Factor graph of an active inference model commonly used for kinematics. Hidden causes of a single level represent a target to reach, while hidden states define the joint angles of the kinematic chain, generating proprioceptive predictions through $\bm{g}_{p}$ and visual predictions through $\bm{g}_{v}$. Forward and inverse kinematics are duplicated, further requiring the likelihood $\bm{g}_{v}$ to be embedded into the dynamics function. (b) Factor graph of an alternative hierarchical model for kinematics. Two different units (left and right blocks) encode information about arm joint angles and target position, respectively generating proprioceptive and visual predictions. The two levels are linked by the likelihood function $\bm{g}_{e}$ performing forward kinematics. Inverse kinematics and goal-directed behavior arise naturally through inference; moreover, both levels can express their own dynamics, affording more advanced control. (c) Generative model of (b).
The visual likelihood function $\bm{g}_{v}$ generates visual predictions for the hand and the target through forward kinematics and identity map, respectively. For goal-directed behavior, first an error between the target and hand positions is generated; then, an inverse kinematic model is embedded directly into the dynamics function, e.g., a Jacobian transpose or a pseudoinverse [70, 71, 72, 53, 32, 81, 79]. In other words, an extrinsic reference frame is inverted to generate an intrinsic state, which is in turn transformed again in the first domain to be compared with visual observations. As a result, forward and inverse kinematics are performed twice, once in the dynamics function and once in the forward and backward passes of visual inference, i.e., when backpropagating the visual prediction error $\bm{\varepsilon}_{o,v}$:
$\partial\bm{g}_{v}^{T}\bm{\varepsilon}_{o,v}=\begin{bmatrix}\bm{J}\\ \bm{1}\end{bmatrix}(\bm{o}_{v}-\begin{bmatrix}\bm{T}(\bm{x})&\bm{v}\end{bmatrix})$ (19)
If the predictions are not temporarily stored, this requires increased computational demand and memory. Additionally, there is an issue regarding biological plausibility: using sensory-level attractors within the dynamics function means that a unit is aware of part of the likelihood prediction – generally assumed to go all down to the sensorium – and its inverse mapping, which are lower-level features. Finally, with the model in Figure 7(a) the agent cannot easily express paths in extrinsic coordinates needed, e.g., for realizing linear or circular motions, or for imposing constraints in both intrinsic and extrinsic domains such as when walking with a glass in hand.
@@
We can instead exploit forward and inverse kinematics of Equation 19 and follow the natural flow of the generative process to avoid duplicated computations, as displayed in Figure 7(b). The alternative generative model is displayed in Figure 7(c). This model relies on two hierarchical levels, where an intrinsic unit (encoding the arm joint angles) is placed at the top and generates predictions through forward kinematics for an extrinsic unit (encoding the Cartesian position of the target) [80]:
$\bm{x}_{e}=\bm{g}_{e}(\bm{x}_{i})+\bm{w}_{o,e}=\bm{T}(\bm{x}_{i})+\bm{w}_{o,e}$ (20)
@@
while the visual likelihood becomes a simple identity map, i.e., $\bm{g}_{v}(\bm{x}_{e})=\bm{x}_{e}$. The goal-directed behavior from the dynamics function of Equation 18 arises naturally via backpropagation of the extrinsic prediction error $\bm{\varepsilon}_{o,e}$:
$\partial\bm{g}_{e}^{T}\bm{\varepsilon}_{o,e}=\bm{J}^{T}(\bm{\mu}_{x,e}-\bm{T}(\bm{\mu}_{x,i}))$ (21)
@@
Having a complete unit that deals with extrinsic information – which is thus not embedded into the hidden causes of the intrinsic unit – allows the agent to specify its dynamics, leading to an efficient decomposition between intrinsic and extrinsic attractors – i.e., $\bm{f}_{i}(\bm{x}_{i},\bm{v}_{i})$ and $\bm{f}_{e}(\bm{x}_{e},\bm{v}_{e})$ – and between proprioceptive and visual observations – as exemplified in the simulations of Figure 8. Note the similarity of Equation 21 with Equations 18 and 19: if in the model of Figure 7(a) we had two different forward and inverse kinematics either for goal-directed behavior or for predicting current observations, in the new model of Figure 7(b) what is compared with the observations already contains a bias toward intentional states, without the need for sensory-level attractors within the dynamics function. The extrinsic prediction error of Equation 21 becomes zero only when the extrinsic hidden states $\bm{x}_{e}$ match the predictions of the intrinsic unit; if such predictions are not met, $\bm{\varepsilon}_{o,e}$ will flow through the hierarchy eventually generating an action. The update rules for the 0th orders of the intrinsic and extrinsic hidden states are the following:
$\dot{\bm{\mu}}_{x,i}$ $=\bm{\mu}_{x,i}^{\prime}-\bm{\Pi}_{\eta,x,i}\bm{\varepsilon}_{\eta,x,i}+\partial_{x_{i}}\bm{g}_{p}^{T}\bm{\Pi}_{o,p}\bm{\varepsilon}_{o,p}+\partial_{x_{i}}\bm{g}_{e}^{T}\bm{\Pi}_{o,e}\bm{\varepsilon}_{o,e}+\partial_{x_{i}i}\bm{f}_{i}^{T}\bm{\Pi}_{x,i}\bm{\varepsilon}_{x,i}$ (22)
$\dot{\bm{\mu}}_{x,e}$ $=\bm{\mu}_{x,e}^{\prime}-\bm{\Pi}_{o,e}\bm{\varepsilon}_{o,e}+\partial_{x_{e}}\bm{g}_{v}^{T}\bm{\Pi}_{o,v}\bm{\varepsilon}_{o,v}+\partial_{x_{e}}\bm{f}_{e}^{T}\bm{\Pi}_{x,e}\bm{\varepsilon}_{x,e}$ (23)
@@
Although the generative model follows the forward flow of optimal control, the relationship between proprioceptive consequences and extrinsic causes peculiar to active inference still holds because the kinematic inversion regards a high-level process that manipulates abstract (intrinsic or extrinsic) representations, and both of them concur to generate low-level proprioceptive states. As Adams and colleagues note, “The key distinction is not about mapping from desired states in an extrinsic (kinematic) frame to an intrinsic (dynamic) frame of reference, but the mapping from desired states (in either frame) to motor commands[38]. Having said this, there is a significant difference between the two models represented in Figure 7, which can be compared to the two supervised learning modes of predictive coding [25]: a forward mode that fixes the latent states to the labels and the observations to the data can generate highly accurate images of digits, while the inverse classification task is more difficult as there is no univocal mapping between labels and data; instead, a backward mode that fixes the latent states to the data and the observations to the labels achieves high performances on classification but falls short when generating images. Based on this, we can interpret the model of Figure 7(a) as a backward mode that would rapidly generate a proper kinematic configuration with the hand at the target, but that would hardly infer from proprioception the hand position needed to plan movements. Conversely, we can interpret the model of Figure 7(b) as a forward mode that would generate with high accuracy the hand position, but that would find it difficult to infer the kinematic configuration needed to actually realize movement.
@@ 3.2 A module for iterative transformations
The model in Figure 7(b) introduced a hierarchical dependency between two (intrinsic and extrinsic) levels, made possible by connecting hidden states. Instead, the typical approach in continuous-time active inference involves connections between hidden states and causes of a level and hidden causes (and not hidden states) of the subordinate level, as shown in Figure 9(a). While this allows one to impose a trajectory for the unit below, specifying fixed setpoints to the 0th-order hidden states is not as straightforward, since the dynamics prediction error generated from the hidden causes would have to travel back to the previous temporal orders. As clear from Figure 7(b), a connection between hidden states is of high utility when designing hierarchical models. In fact – as represented in Figure 9(b) – it is fundamental in defining the initial states of different temporal scales in discrete models, e.g., for pictographic reading [88] or linguistic communication [88]. Similar connections are used in standard PCNs, wherein each neuron of a level computes a combination of neurons of the level above passed to an activation function [25] – as shown in Figure 9(c).
@@
Figure 9: (a) Factor graph of the typical connection between two continuous levels. Hidden states $\bm{x}^{(i)}$ and hidden causes $\bm{v}^{(i)}$ of level $i$ generate – through the likelihood function $\bm{g}^{(i)}$ – the hidden causes $\bm{v}^{(i+1)}$ of level $i+1$. (b) Connections between two discrete levels. Hidden states $\bm{s}_{1}^{(i)}$ of level $i$ generate – through the prior matrix $\bm{D}^{(i)}$ – the hidden states $\bm{s}_{1}^{(i+1)}$ of level $i+1$. We will cover discrete models in the next chapter. (c) Connections between two levels in PCNs. The likelihood function $\bm{g}^{(i)}$ performs a simple combination of neurons, passed to a nonlinear activation function. Standard PCNs only permit representing the causal structure of a system, without modeling internal dynamics. (d) Factor graph of a level with multiple inputs and outputs from independent units, with similar connectivity to temporal predictive coding. The observation $\bm{o}^{(i,j)}$ becomes the 0th-order hidden state $\bm{x}^{(i+1,j)}$ of the level below, while the prior $\bm{\eta}_{x}^{(i,j)}$ becomes the 0th-order hidden state $\bm{x}^{(i-1,j)}$ of the level above. (e) A network of IE modules. An extrinsic $\bm{x}_{e}^{(i-1)}$, along with an intrinsic signal $\bm{x}_{i}^{(i,j)}$ (e.g., angle for rotation or length for translation), is passed to a function $\bm{g}_{e}^{(i,j)}$, generating a new extrinsic signal $\bm{x}_{e}^{(i,j)}$. (f) Generative model of (e).
@@
Recall that, in the previous kinematic model, the prediction $\bm{g}_{e}(\bm{x}_{i})$ of the intrinsic hidden states acted as a prior for the extrinsic hidden states, while the latter acted as an observation for the intrinsic hidden states. Following this example, we use the observation of a level to bias the 0th-order hidden states of the level below directly:
$\bm{\eta}_{x}^{(i+1)}$ $\equiv\bm{g}^{(i)}(\bm{\mu}_{x}^{(i)})$ (24)
$\bm{o}^{(i)}$ $\equiv\bm{x}^{(i+1)}$ (25)
As a result, the observation prediction error $\bm{\varepsilon}_{o}$ and the prior prediction error $\bm{\varepsilon}_{\eta,x}$ of Equation 4 is expressed by the same variable:
$\bm{\varepsilon}_{o}^{(i)}=\bm{\varepsilon}_{\eta,x}^{(i+1)}=\bm{\mu}_{x}^{(i+1)}-\bm{g}^{(i)}(\bm{\mu}_{x}^{(i)})$ (26)
@@
where the hierarchical level is indicated with a superscript and lower levels are denoted by increasing numbers. We can then design a multiple-input and multiple-output system wherein a level imposes and receives priors and observations to several independent units, as in Figure 9(d). The computation of the free energy in Equation 3 remains unchanged, and the update of the hidden states turns into the following:
$\dot{\tilde{\bm{\mu}}}_{x}^{(i,j)}=\begin{bmatrix}\bm{\mu}_{x}^{\prime{(i,j)}}-\bm{\Pi}_{o}^{(i-1)}\bm{\varepsilon}_{o}^{(i-1)}+\sum_{l}\partial_{x^{(i,j)}}\bm{g}^{(i,l)T}\bm{\Pi}_{o}^{(i,l)}\bm{\varepsilon}_{o}^{(i,l)}+\partial_{x^{(i,j)}}\bm{f}^{(i,j)T}\bm{\Pi}^{(i,j)}_{x}\bm{\varepsilon}^{(i,j)}_{x}\\ \\ -\bm{\Pi}^{(i,j)}_{x}\bm{\varepsilon}^{(i,j)}_{x}\end{bmatrix}$ (27)
where the superscript notation $(i,j)$ indicates the $i$th hierarchical level and the $j$th element within the same level. As evident, this is a similar connectivity to hierarchical dynamical models in temporal predictive coding [21, 89]. Here, the forward prediction error $\bm{\varepsilon}_{o}^{(i-1)}$ combines the predictions from the units above and acts as prior (the gradient is absent since they are expressed in the same domain of the unit considered), while the backward prediction errors $\bm{\varepsilon}_{o}^{(i,l)}$ contain the observations from the units below.
@@
What advantages do deep hierarchical models carry compared to a shallow agent? Although the structure of Figure 7(b) affords a more advanced control with respect to the model in Figure 7(a), its uses are still limited to solving simple tasks, e.g., performing operations with the hand. While simultaneous coordination of multiple limbs is possible, it would require complex dynamics functions, with complexity increasing with the number of joints and ramifications of the kinematic chain. Critically, a shallow agent would not be capable of capturing the hierarchical causal relationships inherent to the generative process, allowing one to predict and anticipate the local exchange of forces that would unfold during movement. As mentioned in the Introduction, a deep model is also required if one has to use tools for manipulation tasks. Besides roto-translations in forward kinematics, iterative transformations are also essential in computer vision – where an image can be subject to scaling, shearing, or projection – and, more in general, whenever changing the basis of a coordinate vector.
@@
For these reasons, we can generalize the last model and construct an Intrinsic-Extrinsic (or IE) module [80, 83, 90]. This module is composed of two units and its role is to perform iterative transformations between reference frames. In brief, a unit $\mathcal{U}_{e}^{(i-1)}$ encodes a signal in an extrinsic reference frame, while a second unit $\mathcal{U}_{i}^{(i)}$ contains a generic intrinsic transformation. Applying the latter to the first signal returns a new extrinsic reference frame embedded in a unit $\mathcal{U}_{e}^{(i)}$. Then, we can define a likelihood function $\bm{g}_{e}$ such that:
$\bm{x}_{e}^{(i)}=\bm{g}_{e}^{(i)}(\bm{x}_{i}^{(i)},\bm{x}_{e}^{(i-1)})+\bm{w}_{o,e}^{(i)}=\bm{T}^{(i)}(\bm{x}_{i}^{(i)})^{T}\bm{x}_{e}^{(i-1)}+\bm{w}_{o,e}^{(i)}$ (28)